OpenStudy (anonymous):

solve 3x^2 + 4x = 2 using quadratic formula

OpenStudy (anonymous):

\[-b +- \sqrt{a^2 - 4ac} /2\]

OpenStudy (anonymous):

\[-4 +- \sqrt{16 - (-24)}\]

OpenStudy (anonymous):

4 +- √40 -------- 2a 4 +- 6.32 -------- 6

OpenStudy (callisto):

@jazy Well... quadratic formula is \[x= \frac{-b \pm \sqrt {b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

4 + 6.32 = 10.31 /6 = 1.72

OpenStudy (anonymous):

Thats what I meant:)

OpenStudy (anonymous):

Thank you guys for helping!

OpenStudy (anonymous):

4 - 6.32 = -2.32 /6 = -0.38

OpenStudy (callisto):

a =3 b=4 c=-2 in this case, so, it becomes \[x= \frac{-4 \pm \sqrt {4^2 - 4(3)(-2)}}{2(3)}\]

OpenStudy (anonymous):

Your solutions are (0, 1.72) and (0,-.38) if Im not mistaken.

OpenStudy (anonymous):

@Callisto how can I type formulas on here?

OpenStudy (callisto):

Use ''equation'' fraction -> \frac{x}{y}

OpenStudy (anonymous):

I mean how do you write for example the quad formula? I've tried but idk how:/\[x = (-b +- \sqrt{b^2 - 4ac}) /2\]

OpenStudy (anonymous):

/2a

OpenStudy (callisto):

lol \frac {-b \pm \sqrt{b^{2} -4ac}}{2a}